carmen sold play tickets. each adult ticket cost $4.00 and each students ticket cost $2.00. The total number of tickets sold was 260, and the total income was $700.00. How many of each kind of tickets were sold?
Answer by KnightOwlTutor(293)

X=# of adult tickets
260-X= # of student tickets
$ made by selling adult tickets is $4(X)
$ made by selling student tickets is $(260-X)($2)
The equation is 4X +(2)(260-X) =$700
Use the distributive property
$4X + $520 -$2X=$700
Add $4X -$2X= $2X
$2X+ $520=$700
Subtract $520 from both sides
$2X=$180
X=$90

Therefore the # of adult tickets sold were 90, 260-90=# of student tickets sold=170. To confirm plug in values. $4(90)+ 2(170)=$700
$360 + $340= $700